Cullumber Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $42 throughout the country to loyal alumni of over 3,300 schools. Cullumber’s variable costs are 42% of sales; fixed costs are $116,000 per month.

Assume that variable costs increase to 45% of the current sales price and fixed costs increase by $14,300 per month. If Cullumber were to raise its sales price 11% to cover these new costs, but the number of blankets sold were to drop by 6%, what would be the new annual operating income? (Round sales price to 2 decimal places, e.g. 52.75 and final answer to 0 decimal places, e.g. 5,275.)

139,000 blankets sold by year
Answer #1

New selling price = $42 + 11% = $46.62

New variable cost = $42 X 45% = $18.90

New fixed cost = ($116,000 + $14,300) X 12 = $1,563,600

Number of units sold = 139,000 - 6% = 130,660

New annual operating income =

[(New selling price - New variable cost) X Number of units sold] - New fixed cost

= [($46.62 - $18.90) X 130,660] - $1,563,600

= $2,058,295